Suppose the price of Fanta falls from R18 to R14 per bottle, and the quantity demanded rises from 250 to 400 bottles a day. 

 

Calculate the price elasticity of demand using the point elasticity formula and

Comment on the behaviour of demand for Fanta, based on the price elasticity of demand calculated in 1.1.1 above 


1
Expert's answer
2021-11-08T11:54:03-0500

Using the point elasticity formula, price elasticity of demand


= change in quantity demanded"\div" change in price


="\frac{\delta Q}{\delta P}.\frac{P}{Q}"


="\frac{\ (400-250)}{\ (14-18)}.\frac{14}{250}=-2.1"


Based on the price elasticity of demand calculated above, demand for Fanta will decrease when price increases, and its demand will increase when price decreases.